# Flexible Deep Neural Networks for Partially Linear Survival Data

### v1

@@ 4.3 Asymptotic Results
The proofs of all lemmas and theorems in this section appear in the appendix. We begin by establishing several basic properties of the model. A key step is to understand how the counting–process framework allows us to translate expressions involving at-risk process into simpler quantities expressed directly in terms of the event indicator $\Delta$. The next lemma formalizes this relationship, showing that these two representations are equivalent up to a multiplicative constant. This equivalence will be used repeatedly throughout the analysis.
@@
Under Assumptions A1A2, for any nonnegative, predictable function $f(t,{\bf X},{\bf Z})$ with $\mathbb{E}\!\left[\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right]<\infty$,
@@
$c_{1}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,$, (7)
$c_{3}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$ (8)
In particular, if $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$, then $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–a.s. and $\phi(t,{\bf X},{\bf Z})=0$ for Lebesgue-a.e. $t\in[0,\tau]$ and $P_{{\bf X},{\bf Z}}$–a.e. $({\bf X},{\bf Z})$, i.e. $\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$ for $dt\otimes P_{{\bf X},{\bf Z}}$–almost all $(t,{\bf X},{\bf Z})$.
Lemma 2 shows that the distance $d(\cdot,\cdot)$ is a valid metric on the parameter space: $d(\eta_{1},\eta_{2})=0$ if and only if the induced linear score coincide almost everywhere. In addition, the lemma shows that the event-weighted norm and the corresponding unweighted norms are equivalent up to multiplicative constants. As a result, any bound or convergence statement expressed in terms of the metric $d$ can be translated into an equivalent statement under the usual $L^{2}$ norm and vice versa. This equivalence is used repeatedly throughout the theoretical development.
Next we study the shape of the population loss around $\eta_{o}=(g_{o},\mbox{${\theta}$}_{o})$.
@@
Assume A1A2. For any $\eta\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$,
@@
We begin by quantifying how well the estimator $\widehat{g}$ recovers the unknown function $g_{o}$. The next theorem gives a nonparametric convergence rate for $\widehat{g}$ in the event–weighted $L^{2}_{P_{0}}$ norm that naturally appears in the loss and in the distance function $d$. The rate is expressed in terms of the complexity parameter $\gamma_{n}$ and will later be used to control remainder terms in the expansion of the efficient score.
@@
$\left(\mathbb{E}\!\Bigl[\Delta\bigl\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\Bigr]\right)^{1/2}=O_{p}(\tau_{n})$ (9)
where the expectation is taken with respect to an independent draw $(T,\Delta,{{\bf X}},{{\bf Z}})\sim P_{0}$ from the true distribution, and $\tau_{n}=\gamma_{n}\log^{2}n$. The quantity in (9) is random through its dependence on the training sample used to construct $\widehat{g}$.
The conclusion of Theorem 1 is strengthened by Lemma 2, which shows that $d(\eta,\eta_{0})^{2}$ is equivalent to the unweighted norm $\|\chi_{\eta}-\chi_{\eta_{0}}\|^{2}_{L^{2}(dt\otimes P_{\mathbf{X},\mathbf{Z}})}$. Consequently, the estimator converges in $L^{2}$ over the entire time domain $(0,\tau)$ with respect to Lebesgue measure, and not merely at observed event times or in a way that depends on the censoring pattern.
The next result shows that the rate obtained in Theorem 1 is essentially optimal, up to logarithmic factors. It establishes a minimax lower bound for estimation of $g_{o}$ under the same event–weighted $L^{2}_{P_{0}}$ loss, taken uniformly over the parameter space $\Theta$ and the nuisance function class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
@@
Under A1A2 and A5 there exists $c>0$ such that
$\inf_{\widehat{g}}\sup_{\theta_{o}\in\Theta,\,g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)}\mathbb{E}_{P_{g_{o},\theta_{o}}}\!\left[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\right]\;\;\geq\;c\,\gamma_{n}^{2}\,.$
Next, we turn to the parametric component $\mbox{${\theta}$}_{o}$ and characterize the efficient score and information bound in our semiparametric model. To characterize the nuisance tangent space associated with $g_{o}$, we consider differentiable one-dimensional submodels that pass through $g_{o}$. Let $\mathcal{H}_{g_{o}}$ denote the collection of all functions $g\in L^{2}([0,\tau]\times[0,1]^{r})$ for which there exist a submodel $\{g_{b}:b\in(-1,1)\}\subset\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$ satisfying
@@
$\mathcal{T}_{g_{o}}=\Big\{g\in L^{2}([0,\tau]\times[0,1]^{r}):\exists\,\{g_{b}\}\in\mathcal{H}_{g_{o}}\text{ s.t. }\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|=0\Big\}\,.$
Finally, semiparametric efficiency theory requires a closed linear subspace, so we take the closure of the linear span of $\mathcal{T}{g_{o}}$, denoted $\overline{\mathcal{T}}{g_{o}}$, which forms the nuisance tangent space.
@@
$\mathbb{E}\!\left\{\Delta\,\|{\bf Z}-{\bf g}^{*}(T,{\bf X})\|_{c}^{2}\right\},\quad\|v\|_{c}^{2}=(v_{1}^{2},\ldots,v_{p}^{2})^{\top}\,$,
so ${{\bf g}}^{*}(t,{\bf X})$ projects the covariates $\bf Z$} onto the tangent space of the nuisance function. Finally, under the counting-process representation of the model (andersen1982cox), the counting process $N(t)=\Delta I(T\leq t)$ admits the Doob-Meyer decomposition where the associated martingale is
@@
Under Assumptions A1A2 and A5, the efficient score for $\mbox{${\theta}$}_{o}$ in the semiparametric model (1) is
@@
where ${\bf g}^{*}(t,{\bf X})$ is the projection of $\bf Z$} onto the nuisance tangent space and $M(t)$ is the associated counting–process martingale. The semiparametric information bound for $\mbox{${\theta}$}_{0}$ is
@@
The key technical observation is that, in the Taylor expansion of the population score around $g_{o}$, the first–order derivative in $\widehat{g}-g_{o}$ direction vanishes because the efficient score is orthogonal to the nuisance tangent space. The remaining second–order term is of smaller order, specifically $o_{p}(n^{-1/2})$, provided that $n\gamma_{n}^{4}\to 0$ as $n\to\infty$.
@@
$\delta_{n}\;=\;\max_{1\leq j\leq m_{n}}(t_{j}-t_{j-1})$
denote the mesh size of this grid. Assume that $\delta_{n}=o(\tau_{n}^{2})$ and $\sqrt{n}\,\delta_{n}\;\rightarrow\;0$, as $n\rightarrow\infty$, and that the conditions of Theorems 1 and 4 hold. Then,
@@
[…]

### v2

@@ 4.2 Asymptotic Results
The proofs of all lemmas and theorems in this section appear in Appendix A. We begin by establishing several basic properties of the model. A key step is to understand how the counting–process framework allows us to translate expressions involving the at-risk processes into simpler quantities expressed directly in terms of the event indicator $\Delta$. The next lemma formalizes this relationship, showing that these two representations are equivalent up to a multiplicative constant. This equivalence will be used repeatedly throughout the analysis.
@@
Under A1A2, for any nonnegative, predictable function $f(t,{\bf X},{\bf Z})$ with $\mathbb{E}\!\left[\int_{0}^{\tau}Y(t)\,f(t,{\bf X},{\bf Z})\,dt\right]<\infty$,
@@
$c_{1}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}\!\big\{\phi(T,{\bf X},{\bf Z})^{2}\big\}\;\leq\;c_{2}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,$, (4.2)
$c_{3}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}$ $\;\leq\;\mathbb{E}_{{\bf X},{\bf Z}}\!\Big\{\int_{0}^{\tau}\phi(t,{\bf X},{\bf Z})^{2}\,dt\Big\}\;\leq\;c_{4}\,\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}\,.$ (4.3)
In particular, if $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$, then $\phi(T,{\bf X},{\bf Z})=0$ $\mathbb{P}_{0}$–a.s. and $\phi(t,{\bf X},{\bf Z})=0$ for Lebesgue-a.e. $t\in[0,\tau]$ and $\mathbb{P}_{{\bf X},{\bf Z}}$–a.e. $({\bf X},{\bf Z})$, i.e. $\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})$ for $dt\otimes\mathbb{P}_{{\bf X},{\bf Z}}$–almost all $(t,{\bf X},{\bf Z})$.
Lemma 2 shows that the distance $d(\cdot,\cdot)$ is a valid metric on the parameter space: $d(\eta_{1},\eta_{2})=0$ if and only if the induced linear score coincides almost everywhere. In addition, the lemma shows that the event-weighted norm and the corresponding unweighted norms are equivalent up to multiplicative constants. As a result, any bound or convergence statement expressed in terms of the metric $d$ can be translated into an equivalent statement under the unweighted $L^{2}(dt\otimes P_{{\bf X},{\bf Z}})$ norm and vice versa. This equivalence is used repeatedly throughout the theoretical development.
Next we study the shape of the population loss around $\eta_{o}=(\mbox{${\theta}$}_{o},g_{o})$.
@@
Assume A1A2. For any $\eta\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$,
@@
We begin by quantifying how well the estimator $\widehat{g}$ recovers the unknown function $g_{o}$. The next theorem gives a nonparametric convergence rate for $\widehat{g}$ in the event–weighted $L^{2}_{\mathbb{P}_{0}}$ norm that naturally appears in the loss and in the distance function $d(\eta,\eta_{o})$. The rate is expressed in terms of the complexity parameter $\gamma_{n}$ and will later be used to control remainder terms in the expansion of the efficient score.
@@
$\left(\mathbb{E}\!\Bigl[\Delta\bigl\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\bigr\}^{2}\Bigr]\right)^{1/2}=O_{p}(\tau_{n})$ (4.4)
where the expectation is taken with respect to an independent draw $(T,\Delta,{{\bf X}},{{\bf Z}})\sim\mathbb{P}_{0}$ from the true distribution, and $\tau_{n}=\gamma_{n}\log^{2}n$. The quantity in (4.4) is random through its dependence on the training sample used to construct $\widehat{g}$.
The conclusion of Theorem 1 is strengthened by Lemma 2, which shows that $d(\eta,\eta_{o})^{2}$ is equivalent to the unweighted norm $\|\chi_{\eta}-\chi_{\eta_{0}}\|^{2}_{L^{2}(dt\otimes\mathbb{P}_{\mathbf{X},\mathbf{Z}})}$. Consequently, the estimator converges in $L^{2}$ over the entire time domain $(0,\tau)$ with respect to Lebesgue measure, and not merely at observed event times or in a way that depends on the censoring pattern.
In Theorem 1, both Zhong et al. (2022) and our work rely on Schmidt-Hieber (2020) exclusively for deterministic approximation results of sparse ReLU neural networks over composite Hölder function classes, including the network size and sparsity scaling required to achieve optimal approximation rates. These results are purely analytic and do not involve any distributional assumptions. All stochastic components of the proof, such as control of the estimation error, empirical process bounds, and derivation of convergence rates, are fundamentally different and are derived from the survival-analysis framework. In Zhong et al. (2022), these arguments rely on the Cox partial likelihood and counting-process martingale theory, while in our setting they are based on the full likelihood and properties of the hazard function. Consequently, unlike Schmidt-Hieber (2020), no assumptions on an additive error distribution, including normality, are required or used at any stage of the analysis.
The next result shows that the rate obtained in Theorem 1 is essentially optimal, up to logarithmic factors. It establishes a minimax lower bound for estimation of $g_{o}$ under the same event–weighted $L^{2}_{\mathbb{P}_{0}}$ loss, taken uniformly over the parameter space $\Theta$ and the nuisance function class $\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$.
@@
Under A1A2 and A4 there exists $c>0$ such that
$\inf_{\widehat{g}}\sup_{\theta_{o}\in\Theta,\,g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)}\mathbb{E}_{\mathbb{P}_{g_{o},\theta_{o}}}\!\left[\Delta\{\widehat{g}(T,{\bf X})-g_{o}(T,{\bf X})\}^{2}\right]\;\;\geq\;c\,\gamma_{n}^{2}\,.$
We now shift from nuisance-rate optimality to semiparametric efficiency for the finite-dimensional parameter $\mbox{${\theta}$}_{o}$ and characterize the efficient score and information bound in our semiparametric model. To characterize the nuisance tangent space associated with $g_{o}$, we consider differentiable one-dimensional submodels that pass through $g_{o}$. Let $\mathcal{H}_{g_{o}}$ denote the collection of all functions $g\in L^{2}([0,\tau]\times[0,1]^{r})$ for which there exist a submodel $\{g_{b}:b\in(-1,1)\}\subset\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$ satisfying
@@
$\mathcal{T}_{g_{o}}=\Big\{g\in L^{2}([0,\tau]\times[0,1]^{r}):\exists\,\{g_{b}\}\in\mathcal{H}_{g_{o}}\text{ s.t. }\lim_{b\to 0}\left\|b^{-1}(g_{b}-g_{o})-g\right\|_{L^{2}([0,\tau]\times[0,1]^{r})}=0\Big\}\,.$
@@
$g_{j}^{*}\in\arg\min_{g\in\overline{\mathcal{T}}_{g_{o}}}\mathbb{E}\!\left[\Delta\{Z_{j}-g(T,{\bf X})\}^{2}\right],\qquad j=1,\ldots,p\,.$
Equivalently,
${\bf g}^{*}\in\arg\min_{{\bf g}\in(\overline{\mathcal{T}}_{g_{o}})^{p}}\mathbb{E}\!\left[\Delta\|{\bf Z}-{\bf g}(T,{\bf X})\|_{2}^{2}\right].$
Since $(\overline{\mathcal{T}}_{g_{o}})^{p}$ is a Cartesian product and the criterion separates across coordinates, this minimization is componentwise. Thus, ${\bf g}^{*}(t,{\bf X})$ is the $\Delta$-weighted $L_{2}$-projection of $\bf Z$} onto $(\overline{\mathcal{T}}_{g_{o}})^{p}$. Finally, under the counting-process representation of the model (Andersen and Gill, 1982), the counting process $N(t)=\Delta I(T\leq t)$ has compensator $\int_{0}^{t}Y(s)\exp\{\mbox{${\theta}$}_{o}^{\top}{\bf Z}+g_{o}(s,{\bf X})\}\,ds$ so the associated martingale is
@@
Under Assumptions A1A5, the efficient score for $\mbox{${\theta}$}_{o}$ in the semiparametric model (2.1) is
@@
where ${\bf g}^{*}(t,{\bf X})$ is the $\Delta$-weighted $L_{2}$-projection of $\bf Z$} onto $(\overline{\mathcal{T}}_{g_{o}})^{p}$, and $M(t)$ is the associated counting-process martingale. The semiparametric information bound for $\mbox{${\theta}$}_{0}$ is
@@
The key technical observation is that, in the Taylor expansion of the population score around $g_{o}$, the first–order derivative in $\widehat{g}-g_{o}$ direction vanishes because the efficient score is orthogonal to the nuisance tangent space. The remaining second–order term is $o_{p}(n^{-1/2})$ provided that $n\gamma_{n}^{4}\to 0$ as $n\to\infty$.
@@
$\omega_{n}(\delta)=\sup_{\eta\in\mathcal{N}_{n}}\sup_{{\bf X}\in\mathcal{X}}\sup_{\begin{subarray}{c}s,t\in[0,\tau]\\ |t-s|\leq\delta\end{subarray}}|g(t,{\bf X})-g(s,{\bf X})|\,,\qquad\mathcal{N}_{n}=\{\eta:d(\eta,\eta_{o})\leq M\tau_{n}\}\,$,
for some fixed $M>0$. Assume that $\omega_{n}(\delta_{n})=o(\tau_{n}^{2})$, $\sqrt{n}\,\omega_{n}(\delta_{n})\to 0$, and that the conditions of Theorems 1 and 4 hold. Then,
@@
$\sqrt{n}(\widetilde{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})=\sqrt{n}(\widehat{}\mbox{${\theta}$}-\mbox{${\theta}$}_{o})+o_{p}(1)\;\xrightarrow{\mathcal{D}}\;N\bigl(\mathbf{0},\,I(\mbox{${\theta}$}_{o})^{-1}\bigr)\,$,